A truck traveling at 70 miles/hr consumes 30% more fuel than another that is traveling at 50 miles/hr.If truck traveling at 50 miles/hr has fuel efficiency of 19.5 miles/litre.Then how many miles can the truck traveling at 70 miles/hr travel on 10 litre of fuel.
Answer by mananth(16946)

Truck traveling 50mph consumption =19.5 miles/liter
..
30% more = 19.5 *1.3= 25.35 miles /liter
Truck travelling at 70 mph consumes 25.35 miles/liter
With 10 liters it can go 25.35 *10 =253.5 miles